OpenStudy (anonymous):

Hi everyone.. I am having trouble getting the right answer after a few tries of this double integral. Double integral sin^3u du dv. Evaluate u from o to pi/2 and v from 0 to 2pi.

OpenStudy (anonymous):

\[\int\limits_{o}^{2\pi}\int\limits_{0}^{\pi/2}\sin^3(u)du dv\]

OpenStudy (anonymous):

\[-\int\limits_{0}^{2\pi}\int\limits_{o}^{\pi/2}(1-u^2)du\]

OpenStudy (anonymous):

thats du dv sorry...

OpenStudy (anonymous):

\[-\int\limits_{o}^{2\pi}(u - u^3/3)dv\]

OpenStudy (anonymous):

\[-\int\limits_{o}^{2\pi}(\pi/2 - \pi^3/24)dv\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

oops I see where I went wrong...